# BanditSpec: Adaptive Speculative Decoding via Bandit Algorithms

### v1

@@ C.1 Vanilla Speculative Decoding
For completeness, we present and describe the vanilla speculative decoding algorithm (Leviathan et al., 2023; Chen et al., 2023) in Algorithm 6 in this section.
We introduce some notations first. For any nonnegative function $f:\mathcal{X}\rightarrow\mathbb{R}_{+}$ with $\sum_{x\in\mathcal{X}}f(x)>0$, we define the distribution induced by it as $\text{Norm}(f(\cdot))=f(\cdot)/\sum_{x^{\prime}\in\mathcal{X}}f(x^{\prime})$. The positive part of a function $f$ is denoted as $[f(\cdot)]_{+}=\max\{0,f(\cdot)\}$.
Speculative decoding implements a draft model $Q$ to generate draft tokens and let the target model $P$ verify them in parallel. In practice, the draft model is much smaller than the target model. Thus, the draft token generation (Line 1) can be achieved in a short time. Then we let the target model only forward inference once with these draft tokens as inputs (Line 2). The verification procedures (Lines 4 to 9) are designed to guarantee that the output tokens $x_{1:\tau+1}$ is distributed as the target model $P$. Here, the additional $(\tau+1)$-st accepted token is also called the bonus token.
@@ C.3 Full Description of BanditSpec with UCBSpec
[…]
@@ C.4 Full Description of BanditSpec with EXP3Spec
[…]
@@ Appendix D Proofs of Main Results
[…]
The other results are implied by these two results. For equality, we note that this is already proved by Theorem 1 in Yin et al. (2024), where the equality holds for any specification of the hyperparameters. For the inequality, we note that each implementation of SpecDecSub generates at least one token and at most $L+1$ tokens. Thus, the inequality holds almost surely.
@@
Firstly, the stopping time is now a random, which depends on the generated tokens. This cause trouble when we decompose the reward/regret, as both the rewards and time horizon depend on the history. We tackle this problem in Step 1 by making use of the martingale structure of the rewards sequence.
Secondly, we consider the problem under Assumption 4.1, where the number of accepted tokens can be dependent. This is practical as LLM generates tokens in an autoregressive manner. In contrast, under the commonly seen assumption for the $K$-armed MAB, the rewards are i.i.d. and can be regarded as they have been sampled before the algorithm starts (see Chapter 4 in Lattimore & Szepesvári (2020)). Thus, Chernoff-Hoeffding bound (Lemma F.2) can be directly applied, which cannot be used under Assumption 4.1. We solve this problem in Step 2, by adopting the so-called self-normalized confidence bounds (Abbasi-yadkori et al., 2011).
@@
By the property of speculative decoding in (3) and (30), for any algorithm $\mathtt{ALG}$,  [^3] We clarify that only the tokens up to the $\mathrm{EOS}$ token will be appended to the prefix token sequence in practice. Therefore, the actual number of accepted tokens is between $[1,Y_{I_{\mathrm{ST}(\mathtt{ALG})},\mathrm{ST}(\mathtt{ALG})}]$. While this mismatch can be solved, using $Y_{I_{\mathrm{ST}(\mathtt{ALG})},\mathrm{ST}(\mathtt{ALG})}$ as the number of accepted tokens in the last round will introduce an error of at most $L$ to the token sequence length $\mathrm{len}(\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})})$, which is an error of at most $1$ to the stopping time $\mathrm{ST}(\mathtt{ALG})$. This error is negligible compared to the other values in the stopping time regret. Therefore, we assume the $\mathrm{EOS}$ token only appears at the end of accepted tokens for the sake of simplicity.
@@
The standard regret analysis adopts Wald’s equation to decompose the expected cumulative regret, or equivalently the stopping time. However, as both the stopping time $\mathrm{ST}(\mathtt{ALG})$ and $Y_{I_{t},t}$ depends on the history under our problem setup, Wald’s equation fails. We propose a new and general approach to decompose the reward.
@@
$\mathcal{E}_{t}:=\big{\{}\hat{\mu}_{i,t}\in[\mu_{i}-\mathrm{cr}_{i,t},\mu_{i}+\mathrm{cr}_{i,t}],\forall i\in[K]\mbox{ at round }t\big{\}}.$ (31)
@@
where $\delta$ is a confidence parameter that will be specified later. We remark that Lemma F.3 from Abbasi-yadkori et al. (2011) adopts a self-normalized concentration bound for the martingale sequence, which generalizes the standard i.i.d. reward assumption in the $K$-armed MAB problem.
@@
[…]
@@
Our analysis of Algorithm 10 is novel compared to the standard analysis of EXP3 algorithm (Auer et al., 2002a; Lattimore & Szepesvári, 2020). It requires more technical manipulations due to the fact that the termination times of the baseline algorithm $\mathtt{ALG}_{i}$ and $\mathtt{ALG}$ are different and random, and that our goal is to minimize the stopping time regret (12).
For theoretical analysis, we regard Algorithm 10 as an instantiation of the Follow-the-Regularized-Leader (FTRL) algorithm (Gordon, 1999; Lattimore & Szepesvári, 2020).
@@
  • Upper bound each term in the decomposed regret: we upper bound each term in the decomposed regret. The main difficulty is to deal with the difference in the time scales $\mathrm{ST}(\mathtt{ALG})$ and $\mathrm{ST}(\mathtt{ALG}_{i})$ and the randomness in $\mathrm{ST}(\mathtt{ALG})$. Specifically, because both the loss vectors $\{\widehat{Z}_{t}\}_{t\in\mathbb{N}}$ and the stopping time $\mathrm{ST}(\mathtt{ALG})$ are random, taking expectation of the cumulative loss within $[\mathrm{ST}(\mathtt{ALG}_{i}),\mathrm{ST}(\mathtt{ALG})]$ is non-trivial. We devise Lemma D.1 and Lemma D.2 to deal with this problem.
@@
Under our problem setup, the difference lies at the target of the algorithm. Instead of minimizing the corresponding regret
@@
where $e_{i}\in\mathbb{R}^{K}$ is an one-hot vector with the $i^{th}$ coordinate being $1$, and the expectation is taken w.r.t. the internal randomness within $\mathtt{ALG}$ and $\widehat{Z}_{t}$. We highlight again that $\mathrm{ST}(\mathtt{ALG})$ is random whereas $\mathrm{ST}(\mathtt{ALG}_{i})$ is fixed under the greedy decoding strategy.
@@
Based on the previous two steps, we now upper bound $\widetilde{\mathrm{Reg}}(\mathtt{ALG})$. This notion of regret distinguishes from the standard regret analysis due to the difference in the time scales $\mathrm{ST}(\mathtt{ALG})$ and $\mathrm{ST}(\mathtt{ALG}_{i})$.
@@
In this step, we upper bound each term in (64). We comment that ()($\Box$)( □ ) and ()($\P$)( ¶ ) require us to attend to the randomness in $\mathrm{ST}(\mathtt{ALG})$ and the different time indeces. This issue will not be encountered in the conventional scenario where $\mathrm{ST}(\mathtt{ALG})=\mathrm{ST}(\mathtt{ALG}_{i})$.
$\bullet$ Upper bound ()($\Box$)( □ ): we will show that $\mathbb{E}[(\Box)]\leq\mathbb{E}[K\sum_{t=1}^{\mathrm{ST}(\mathtt{ALG})}\eta_{t}/2]$ almost surely.
@@
$\bullet$ Upper bound ()($\diamond$)( ⋄ ): Because $F(x)$ is non-negative on $\bm{\Delta}_{[K]}$ and $F(e_{i})=\log K$, hence,
@@
$\bullet$ Upper bound ()($\dagger$)( † ): Recall that $F(x)=\sum_{i=1}^{K}(x_{i}\log x_{i}-x_{i})+\log K+1$, so $F(x)$ is non-negative for any $x\in\bm{\Delta}_{[K]}$. Additionally, $\eta_{t}=\sqrt{\log K/(K\cdot t)},t\in[\mathrm{ST}(\mathtt{ALG})]$ is a decreasing sequence. Therefore,
@@
$\bullet$ Upper bound ()($\ddagger$)( ‡ ): Since $p_{\mathrm{ST}(\mathtt{ALG})+1}$ is the minimizer of $\Phi_{\mathrm{ST}(\mathtt{ALG})+1}(p)$, we have $(\ddagger)\leq 0$.
$\bullet$ Upper bound ()($\P$)( ¶ ): Under Assumption 5.2, $\mathrm{ST}(\mathtt{ALG})\geq\mathrm{ST}(\mathtt{ALG}_{1})$ almost surely. We further prove that $\mathbb{E}[(\P)]\leq\mathbb{E}[\mathrm{ST}(\mathtt{ALG})]-\mathrm{ST}(\mathtt{ALG}_{i})$ for $i=i^{*}$, which we summarize in the following lemma with proof postponed to Appendix E.1.
@@
Step 2: Establishment for the lower bound of $\mathbb{E}_{\nu}[n_{i,\mathrm{ST}(\mathtt{ALG})}]$. Given algorithm $\mathtt{ALG}$, bandit configuration $\nu\in\Lambda$, prompt $\mathrm{pt}\in(\mathrm{pt}^{m})_{m=1}^{\infty}$ and any $\varepsilon>0$, construct $K-1$ alternative bandit configurations $\nu_{i}=(P,\mathcal{S}_{i}=\{S_{i,j}\}_{j=1}^{K},L)$ for $i\neq i^{*}$ with
@@
where the last equality follows from the definition of consistent algorithm. Because $\varepsilon>0$ is arbitrarily chosen, by sending $\varepsilon\to 0$, we obtain the lower bound for $\mathbb{E}_{\nu}[n_{i,\mathrm{ST}(\mathtt{ALG})}]$
@@ Appendix F Supporting Lemmas
Lemma F.1 (Doob’s optional stopping, Theorem 3.8 in Lattimore & Szepesvári (2020)).
@@
Lemma F.2 (Chernoff-Hoeffding bound, Fact 1 in Auer et al. (2002b)).
@@
$\mathbb{P}[S_{n}\geq n\mu+a]\leq\exp\left(-\frac{2a^{2}}{n}\right)\quad\mbox{ and }\quad\mathbb{P}[S_{n}\geq n\mu-a]\leq\exp\left(-\frac{2a^{2}}{n}\right)$ (139)
Lemma F.3 (Confidence Intervals, Lemma 6 in Abbasi-yadkori et al. (2011)).
@@
Lemma F.4 (Lemma 8 in Antos et al. (2010)).
@@
Lemma F.5 (Exercise 3.7 in Lattimore & Szepesvári (2020)).
@@
Lemma F.6 (Bretagnolle-Huber inequality, Theorem 14.2 in Lattimore & Szepesvári (2020)).
@@
Lemma F.7 (Pinsker’s inequality, Equation (14.12) in Lattimore & Szepesvári (2020)).
@@
Lemma F.8 (Theorem 26.13 in Lattimore & Szepesvári (2020)).
@@ G.1 Additional Experimental Details
For the experiments stated in Section 6.1, we report the memory utilization in this section. As Ealge-2 (Li et al., 2024b) is one of the best speculative decoding methods, we adopt it as the baseline. The Normalized Memory (NM) and Normalized Memory Bandwidth (NMB) are presented in Table 2. The result shows that the proposed methods do not incur additional memory consumption compared to the baseline method.
@@
  • All the draft models share the same verifier model, which is the target model (LLaMA3-8B-Instruct (Dubey et al., 2024) and Qwen2-7B-Instruct (Yang et al., 2024) in our experiments). So that the storage of the verifier does not increase the GPU memory.
@@
We note that the size of SpecBench is not large enough, i.e., the number of arms pulls is not large, to derive a statistically sound result. We enable Mixture-of-Agent (Wang et al., 2024) on the prompts whose responses are shorter than $100$ tokens to increase the number of arm pulls.

### v2

@@ C.1 Vanilla Speculative Decoding
For completeness, we present and describe the vanilla speculative decoding algorithm (leviathan2023fast; chen2023accelerating) in Algorithm 6 in this section.
We introduce some notations first. For any nonnegative function $f:\mathcal{X}\rightarrow\mathbb{R}_{+}$ with $\sum_{x\in\mathcal{X}}f(x)>0$, we define the distribution induced by it as $\text{Norm}(f(\cdot))=f(\cdot)/\sum_{x^{\prime}\in\mathcal{X}}f(x^{\prime})$. The positive part of a function $f$ is denoted as $f(\cdot)]_{+}=\max\{0,f(\cdot)\}$.
Speculative decoding implements a draft model $Q$ to generate draft tokens and let the target model $P$ verify them in parallel. In practice, the draft model is much smaller than the target model. Thus, the draft token generation (Line 1) can be achieved in a short time. Then we let the target model only forward inference once with these draft tokens as inputs (Line 2). The verification procedures (Lines 4 to 9) are designed to guarantee that the output tokens $x_{1:\tau+1}$ are distributed as the target model $P$. Here, the additional $(\tau+1)$-st accepted token is also called the bonus token.
@@ C.3 Full Description of BanditSpec with UCBSpec
Algorithm 8 $\textsc{BanditSpec}(\textsc{UCBSpec})$ (Full version of UCBSpec)
@@ C.4 Full Description of BanditSpec with EXP3Spec
We provide the full description of $\textsc{BanditSpec}(\mathtt{ALG})$ with $\mathtt{ALG}=\textsc{EXP3Spec}$ in Algorithm 9.
Algorithm 9 $\textsc{BanditSpec}(\textsc{EXP3Spec})$ (Full version of EXP3Spec)
@@ Appendix D Proofs of Main Results
$\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})}\overset{d}{=}\mathrm{pt}_{\tau_{\mathrm{c}}},\text{ and }\frac{\mathrm{len}(\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})})}{L+1}\leq\mathrm{ST}(\mathtt{ALG})\leq\mathrm{len}(\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})}),\quad a.s.$
The other results are implied by these two results. For equality, we note that this is already proved by Theorem 1 in yin2024theoretical, where the equality holds for any specification of the hyperparameters. For the inequality, we note that each implementation of SpecDecSub generates at least one token and at most $L+1$ tokens. Thus, the inequality holds almost surely.
@@
Firstly, the stopping time is now a random variable, which depends on the generated tokens. This causes trouble when we decompose the reward/regret, as both the rewards andthe time horizon depend on the history. We tackle this problem in Step 1 by making use of the martingale structure of the rewards sequence.
Secondly, we consider the problem under Assumption 4.1, where the number of accepted tokens can be dependent. This is practical as LLM generates tokens in an autoregressive manner. In contrast, under the commonly seen assumption for the $K$-armed MAB, the rewards are i.i.d. and can be regarded as if they had been sampled before the algorithm starts (see Chapter 4 in lattimore2020bandit). Thus, Chernoff-Hoeffding bound (Lemma F.2) can be directly applied, which cannot be used under Assumption 4.1. We solve this problem in Step 2, by adopting the so-called self-normalized confidence bounds (Abbasi2011improved).
@@
By the property of speculative decoding in (3) and (30), for any algorithm $\mathtt{ALG}$,  [^3] We clarify that only the tokens up to the $\mathrm{EOS}$ token will be appended to the prefix token sequence in practice. Therefore, the actual number of accepted tokens is between $1,Y_{I_{\mathrm{ST}(\mathtt{ALG})},\mathrm{ST}(\mathtt{ALG})}]$. While this mismatch can be solved, using $Y_{I_{\mathrm{ST}(\mathtt{ALG})},\mathrm{ST}(\mathtt{ALG})}$ as the number of accepted tokens in the last round will introduce an error of at most $L$ to the token sequence length $\mathrm{len}(\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})})$, which is an error of at most $1$ to the stopping time $\mathrm{ST}(\mathtt{ALG})$. This error is negligible compared to the other values in the stopping time regret. Therefore, we assume the $\mathrm{EOS}$ token only appears at the end of accepted tokens for the sake of simplicity.
@@
The standard regret analysis adopts Wald’s equation to decompose the expected cumulative regret, or equivalently, the stopping time. However, as both the stopping time $\mathrm{ST}(\mathtt{ALG})$ and $Y_{I_{t},t}$ depend on the history under our problem setup, Wald’s equation fails. We propose a new and general approach to decomposing the reward.
@@
[…]
@@
where $\delta$ is a confidence parameter that will be specified later. We remark that Lemma F.3 from Abbasi2011improved adopts a self-normalized concentration bound for the martingale sequence, which generalizes the standard i.i.d. reward assumption in the $K$-armed MAB problem.
@@
Fix any $b\in[K]$, the baseline algorithm is set to be $\mathtt{ALG}_{b}$, i.e., Algorithm 3 implements Line 4 of Algorithm 3 with hyperparameter $S_{b}$ only. Let $\mathtt{ALG}=\textsc{EXP3Spec}$. we assume the $\textsc{BanditSpec}(\mathtt{ALG})$ repeats the while loop in Algorithm 3 until $\max\{\mathrm{ST}(\mathtt{ALG}),\mathrm{ST}(\mathtt{ALG}_{b})\}$. To avoid any confusion, we restate the algorithm for the purpose of analysis in Algorithm 10. Algorithm 10 takes $\mathtt{ALG}$ and $\mathtt{ALG}_{b}$ as an input and stops until $\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG})}$ and $\mathrm{pt}_{\mathrm{ST}(\mathtt{ALG}_{b})}$ are generated. The two token sequences up to $\mathrm{EOS}$ token are output at the end of the algorithm.
Algorithm 10 $\textsc{BanditSpec}(\textsc{EXP3Spec})$ (For analysis purpose)
@@
Our analysis of Algorithm 10 is novel compared to the standard analysis of EXP3 algorithm (auer2002nonstochastic; lattimore2020bandit). It requires more technical manipulations due to the fact that the termination times of the baseline algorithm $\mathtt{ALG}_{i}$ and $\mathtt{ALG}$ are different and random, and that our goal is to minimize the stopping time regret (12).
For theoretical analysis, we regard Algorithm 10 as an instantiation of the Follow-the-Regularized-Leader (FTRL) algorithm (gordon1999regret; lattimore2020bandit).
@@
  • Upper bound each term in the decomposed regret: we upper bound each term in the decomposed regret. The main difficulty is to deal with the difference in the time scales $\mathrm{ST}(\mathtt{ALG})$ and $\mathrm{ST}(\mathtt{ALG}_{i})$ and the randomness in $\mathrm{ST}(\mathtt{ALG})$. Specifically, because both the loss vectors $\{\widehat{Z}_{t}\}_{t\in\mathbb{N}}$ and the stopping time $\mathrm{ST}(\mathtt{ALG})$ are random, taking expectation of the cumulative loss within $\mathrm{ST}(\mathtt{ALG}_{i}),\mathrm{ST}(\mathtt{ALG})]$ is non-trivial. We devise Lemma D.1 and Lemma D.2 to deal with this problem.
@@
Under our problem setup, the difference lies in the target of the algorithm. Instead of minimizing the corresponding regret
@@
where $e_{i}\in\mathbb{R}^{K}$ is an one-hot vector with the $i^{th}$ coordinate being $1$, and the expectation is taken w.r.t. the internal randomness within $\mathtt{ALG}$ and $\widehat{Z}_{t}$. We highlight again that $\mathrm{ST}(\mathtt{ALG})$ is random, whereas $\mathrm{ST}(\mathtt{ALG}_{i})$ is fixed under the greedy decoding strategy.
@@
Based on the previous two steps, we now upper bound $\widetilde{\mathrm{Reg}}(\mathtt{ALG})$. This notion of regret distinguishes itself from the standard regret analysis due to the difference in the time scales $\mathrm{ST}(\mathtt{ALG})$ and $\mathrm{ST}(\mathtt{ALG}_{i})$.
@@
In this step, we upper bound each term in (64). We comment that $(\Box)$ and $(\P)$ require us to attend to the randomness in $\mathrm{ST}(\mathtt{ALG})$ and the different time indeces. This issue will not be encountered in the conventional scenario where $\mathrm{ST}(\mathtt{ALG})=\mathrm{ST}(\mathtt{ALG}_{i})$.
$\bullet$ Upper bound $(\Box)$: we will show that $\mathbb{E}[(\Box)]\leq\mathbb{E}[K\sum_{t=1}^{\mathrm{ST}(\mathtt{ALG})}\eta_{t}/2]$ almost surely.
@@
$\bullet$ Upper bound $(\diamond)$: Because $F(x)$ is non-negative on $\bm{\Delta}_{[K]}$ and $F(e_{i})=\log K$, hence,
@@
$\bullet$ Upper bound $(\dagger)$: Recall that $F(x)=\sum_{i=1}^{K}(x_{i}\log x_{i}-x_{i})+\log K+1$, so $F(x)$ is non-negative for any $x\in\bm{\Delta}_{[K]}$. Additionally, $\eta_{t}=\sqrt{\log K/(K\cdot t)},t\in[\mathrm{ST}(\mathtt{ALG})]$ is a decreasing sequence. Therefore,
@@
$\bullet$ Upper bound $(\ddagger)$: Since $p_{\mathrm{ST}(\mathtt{ALG})+1}$ is the minimizer of $\Phi_{\mathrm{ST}(\mathtt{ALG})+1}(p)$, we have $(\ddagger)\leq 0$.
$\bullet$ Upper bound $(\P)$: Under Assumption 5.2, $\mathrm{ST}(\mathtt{ALG})\geq\mathrm{ST}(\mathtt{ALG}_{1})$ almost surely. We further prove that $\mathbb{E}[(\P)]\leq\mathbb{E}[\mathrm{ST}(\mathtt{ALG})]-\mathrm{ST}(\mathtt{ALG}_{i})$ for $i=i^{*}$, which we summarize in the following lemma with proof postponed to Appendix E.1.
@@
Step 2: Establishment for the lower bound of $\mathbb{E}_{\nu}[n_{i,\mathrm{ST}(\mathtt{ALG})}]$. Given algorithm $\mathtt{ALG}$, bandit configuration $\nu\in\Lambda$, prompt $\mathrm{pt}\in(\mathrm{pt}^{m})_{m=1}^{\infty}$ and any $\varepsilon>0$, construct K1K-1 alternative bandit configurations $\nu_{i}=(P,\mathcal{S}_{i}=\{S_{i,j}\}_{j=1}^{K},L)$ for $i\neq i^{*}$ with
@@
where the last equality follows from the definition of a consistent algorithm. Because $\varepsilon>0$ is arbitrarily chosen, by sending $\varepsilon\to 0$, we obtain the lower bound for $\mathbb{E}_{\nu}[n_{i,\mathrm{ST}(\mathtt{ALG})}]$
@@ Appendix F Supporting Lemmas
Lemma F.1 (Doob’s optional stopping, Theorem 3.8 in lattimore2020bandit).
@@
Lemma F.2 (Chernoff-Hoeffding bound, Fact 1 in auer2002finite).
@@
[…]
Lemma F.3 (Confidence Intervals, Lemma 6 in Abbasi2011improved).
@@
Lemma F.4 (Lemma 8 in antos2010active).
@@
Lemma F.5 (Exercise 3.7 in lattimore2020bandit).
@@
Lemma F.6 (Bretagnolle-Huber inequality, Theorem 14.2 in lattimore2020bandit).
@@
Lemma F.7 (Pinsker’s inequality, Equation (14.12) in lattimore2020bandit).
@@
Lemma F.8 (Theorem 26.13 in lattimore2020bandit).
@@ G.1 Additional Experimental Details
For the experiments stated in Section 6.1, we report the memory utilization in this section. As Ealge-2 (li2024eagle2) is one of the best speculative decoding methods, we adopt it as the baseline. The Normalized Memory (NM) and Normalized Memory Bandwidth (NMB) are presented in Table 2. The result shows that the proposed methods do not incur additional memory consumption compared to the baseline method.
@@
  • All the draft models share the same verifier model, which is the target model (LLaMA3-8B-Instruct (grattafiori2024llama3herdmodels) and Qwen2-7B-Instruct (yang2024qwen2technicalreport) in our experiments). So that the storage of the verifier does not increase the GPU memory.
@@
We note that the size of SpecBench is not large enough, i.e., the number of arm pulls is not large, to derive a statistically sound result. We enable Mixture-of-Agent (wang2024mixture) on the prompts whose responses are shorter than $100$ tokens to increase the number of arm pulls.